OpenStudy (anonymous):

If you have 2 grams of ammonia and 4 grams of oxygen, how do you write the equation to know your reaction? I know how to balance it once its written, but how do you write the addition of the elements to get the reaction? Been years and I am trying to help my nephew.

OpenStudy (anonymous):

well if you have ammonia at stp then ammonia is in gas state and oxygen is too so: NH3 + O2 -> NO + H2O now you need to balance it, if you have further questions feel free to ask...

OpenStudy (anonymous):

According to this example I have, pertaining to my question, the balanced equation is 4 NH3(g) + 5 O2(g)4 NO(g) + 6 H2O(g), as I said, I know how to balance them once they are written, I have just forgotten how to get from "2 grams of ammonia and 4 grams of oxygen" to "4 NH3(g) + 5 O2(g)4 NO(g) + 6 H2O(g)" Maybe its too much to ask on here, I was just trying to help him out and it seems I know how to balance, but not how to get there lol.

OpenStudy (anonymous):

i see what you are trying to say but any data on reactants and products as mass etc. is not relevant for writing the chemical reaction equation since all reactants react in certain ratio to give certain ratio of products... now i suppose what you need to calculate is what amount of NO is created if you have 2 g of NH3 and 4 g of O2 am i right? if so you just recalculate mass in grams to moles and multiply by reaction ratio which in this case would be: n(NH3) = 2g/ M(NH3) n(O2) = 4g/ M(O2) and when you see which one is limiting reactant you calculate mass or what ever of NO with that amount, for this example without calculation lets say NH3 is limiting reactant: n(NO)/n(NH3)= 4/4 = 1 so: n(NO)=n(NH3)

OpenStudy (anonymous):

Maybe I am not posing this question right, because at this time, I really want to know. If I have 2 grams of ammonia and 4 grams of oxygen and I am mixing them together, what would be the formula "before" you go through the rigamarole of balancing it? i.e.( 2grams of NH3 + 4 grams of O2 >> _NO? + _H2O) What steps get me to the reactant answer? N*O + H3*O >> NO + H2O........then you balance it etc?

OpenStudy (anonymous):

ok so as i said before that mass of ammonia and oxygen is irelevant, it doesnt have any influence on writing balanced chemical equation! the equation is as follows: 2 NH3 + 5/2 O2 -> 2NO + 3H2O and that chemical equation is as it is, reactant mass which is brought to the system or which you get with that reaction is irelevant to the ratio in which reactants react to give products!

OpenStudy (anonymous):

Thanks, but you are not answering my question at all. Perhaps someone else will get what I am asking.